# Learning Massive-scale Partial Correlation Networks in Clinical Multi-omics Studies with HP-ACCORD

### v3

@@ 6.1 HP-ACCORD to delineate complex mechanisms in gene expression regulation
In this section, we show that HP-ACCORD enables sophisticated biological inference using ultrahigh-dimensional multi-omic data through the example of the LIver Hepatocellular Carcinoma (LIHC) cohort ($n$=365) of TCGA. We used the LIHC dataset consisting of the expression data of 15,598 protein-coding genes (mRNAs), the expression data of 364 microRNAs (miRNAs) and DNA methylation levels ($\beta$-values) at 269,396 CpG islands located within the upstream regulatory regions of genes. The methylation data contained a subset of variables provided in the original HumanMethylation450k BeadChip array data from the TCGA data portal [1], where we included the probes in genomic regulatory regions or gene bodies only.
@@ 6.2 Network structure of epigenomic and transcriptomic data
Since the DNA methylation probes outnumbered both mRNA- and miRNA expression variables in the input data, the majority of the selected edges were intra-modality correlations between methylation probes ($>$99%), which showed predominantly positive partial correlations (98%) at all ranges of genomic distances. The second most frequent edges were between mRNAs and between miRNAs, where the vast majority showed positive partial correlations (96%), and the correlations were close to zero for the pairs associated with negative partial correlations. This observation clearly reaffirms that physically proximal protein coding genes and non-coding RNAs are often co-transcribed [34, 37]. Among the small number of non-zero inter-modality correlations, the most pronounced category was that of edges between DNA methylation probes and mRNA expression (total transcript level per gene); the signs of the partial correlations were evenly split between positive and negative (Figure 4A).

### v4

@@ 6.1 HP-ACCORD to delineate complex mechanisms in gene expression regulation
In this section, we show that HP-ACCORD enables sophisticated biological inference using ultrahigh-dimensional multi-omic data through the example of the LIver Hepatocellular Carcinoma (LIHC) cohort (nn=365) of TCGA. We used the LIHC dataset consisting of the expression data of 15,598 protein-coding genes (mRNAs), the expression data of 364 microRNAs (miRNAs) and DNA methylation levels ($\beta$-values) at 269,396 CpG islands located within the upstream regulatory regions of genes. The methylation data contained a subset of variables provided in the original HumanMethylation450k BeadChip array data from the TCGA data portal [1], where we included the probes in genomic regulatory regions or gene bodies only.
@@ 6.2 Network structure of epigenomic and transcriptomic data
Since the DNA methylation probes outnumbered both mRNA- and miRNA expression variables in the input data, the majority of the selected edges were intra-modality correlations between methylation probes (>>99%), which showed predominantly positive partial correlations (98%) at all ranges of genomic distances. The second most frequent edges were between mRNAs and between miRNAs, where the vast majority showed positive partial correlations (96%), and the correlations were close to zero for the pairs associated with negative partial correlations. This observation clearly reaffirms that physically proximal protein coding genes and non-coding RNAs are often co-transcribed [34, 37]. Among the small number of non-zero inter-modality correlations, the most pronounced category was that of edges between DNA methylation probes and mRNA expression (total transcript level per gene); the signs of the partial correlations were evenly split between positive and negative (Figure 4A).